OpenStudy (anonymous):

What is the solution to the rational equation (3)/(x+3)=(2)/(x-2)? Can anyone solve this problem? Any Help would be greatly appreciated!!

OpenStudy (anonymous):

start by "cross multiplying" to get \[3(x-2)=2(x+3)\] and solve that one the usual way distribute first

OpenStudy (anonymous):

Okay i got 3x-6=2x+6 @satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

subtract \(2x\) from both sides , add \(6\) to both sides and you will be done

OpenStudy (anonymous):

is it x=0 @satellite73

OpenStudy (anonymous):

oh its x=12 duh @satellite73